You invest $5,700 for 18 months at an APR of 3.7%. What is the Future Value of your investment?
Answer #1

suppose it compounded daily then the future value of investment = principal*(1+r/n)^n

where r - annual interest rate

n - no, of times componded

future value of invetment = 5700*(1+0.037/360)^540

= 5700*(1.000103)^540

= 5700*1.057066

= 6025.276

and if it compounded monthly

then Future value of investment = 5700*(1+0.037/12)^18

= 5700*(1.003083)^18

= 5700*1.056979

= 6024.779

Please comment in case of any clarification required.